OpenStudy (kabase1234):

I'll fan and medal Find the sum of the arithmetic sequence. -1, 2, 5, 8, 11, 14, 17

OpenStudy (mayankdevnani):

what do you think? any ideas ?

OpenStudy (kabase1234):

No this is new to me

OpenStudy (kabase1234):

@mayankdevnani

OpenStudy (mayankdevnani):

There are two method to solve this :- 1) Find sum using arithmetic progression formula 2) Simply add them

OpenStudy (kabase1234):

I can just add them all together?

OpenStudy (mayankdevnani):

yeah! this is our second method

Atsie (atsie):

Thats what I was going to say was add them all together but I didn't know if there was some catch. :P

OpenStudy (kabase1234):

57?

OpenStudy (kabase1234):

56

OpenStudy (mayankdevnani):

yeah! 56 is correct

OpenStudy (mayankdevnani):

wanna solve it by first method ?

OpenStudy (kabase1234):

How do you do that?

OpenStudy (mayankdevnani):

do you know the formula of summation of arithmetic progression ?

OpenStudy (kabase1234):

No I don't

OpenStudy (mayankdevnani):

If you have given an arithmetic progression with its 1st term and last term,use this formula :- \[\large \bf S_n=\frac{n}{2}(a+l)\] where n=number of terms a=first term l=last term

OpenStudy (mayankdevnani):

given equation :- -1, 2, 5, 8, 11, 14, 17 where a=-1 l=17 n=7 \[\large \bf S_7=\frac{7}{2}(-1+17)=\frac{7}{2} \times 16=8 \times 7=56\]

OpenStudy (mayankdevnani):

you'll get same answer in each case

OpenStudy (kabase1234):

That makes sense! Thank you! Can you help me with one more question??

OpenStudy (mayankdevnani):

sure :D

OpenStudy (kabase1234):

I'll open a new question

OpenStudy (mayankdevnani):

yea sure